OpenStudy (anonymous):

will medal and fan Julie was dividing 13.8 by 0.1. Which statement about the quotient is true? A. It will be greater than 13.8. B. It will be less than 13.8. C. It will be equal to 13.8.

OpenStudy (mathmale):

Why not try this division? Divide 13.8 by 0.1. What result do you obtain?

OpenStudy (anonymous):

i just dont understand on how to do the problem

OpenStudy (anonymous):

@ShiBuringa

OpenStudy (michele_laino):

hint: we can write this: \[\Large \frac{{13.8}}{{0.1}} = \frac{{\left( {\frac{{138}}{{10}}} \right)}}{{\left( {\frac{1}{{10}}} \right)}} = \frac{{138}}{{10}} \times \frac{{10}}{1} = ...?\]

OpenStudy (anonymous):

did the problem and i think it is c.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (mathmale):

I think Michele is waiting for you to share your work. Show us how you "did the problem."

OpenStudy (michele_laino):

please write the result

OpenStudy (michele_laino):

since option C, is a wrong option @amber_nicole

OpenStudy (anonymous):

okay i figured it out its A

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank you!!!!

OpenStudy (michele_laino):

:)